25-20(32+15)-21=x + 2
find x
please let me know the answer quickly
 

Grade:8

1 Answers

Umakant biswal
7 years ago
dear utkarsh 
by solving the equation , we get 
25-640-300-21 = x+2 
25-940-2 = x+2 
+25-942 = x+2 
so, x+2 = -917 
x = -917-2 
x= – 919 
HOPE IT CLEARS 
REGARDS